OpenStudy (anonymous):

If f(x)=4x, g(x)=5x-3, and h(x)=square root of x, find (f o g o h)(4) I keep getting different answers every time I do it, please help!

OpenStudy (anteater):



OpenStudy (anteater):

So you can think of these as "nested" functions, in a sense. Start by doing h(4) which would give you 2. Then do g(2) which gives you 7. And finally f(7) gives 28.

OpenStudy (anonymous):

@anteater Thank you so much! I see what I did wrong now.

OpenStudy (larseighner):

I hate the little circle notation because I like to say work from the inside out; But if you have the little circle notation, work right to left ( @anteater ) has the right answer. But if you don't follow it, start from the right. find h(4) which is the square root of 4, which is 2. Now move left one notch and find g(2). That 2 is your previous result. So 5(2) - 3 = 10 - 3 = 7. Now move one notch to the left and find f(7). Seven was your previous result. f(7) = 4(7) = 28. It is easy to get lost in composite functions, but if you go at it one step at a time, you can often solve simple problems like this one "by inspection."

OpenStudy (anteater):

Welcome!